# DeepTalk: Towards Seamless and Smart Speech Interaction with Adaptive Modality-Specific MoE

### v2

@@ Abstract
Native multimodal large language models (MLLMs) restructure a single large language model (LLM) into a spoken language model (SLM) capable of both speech and text generation. Compared to modular and aligned MLLMs, native MLLMs preserve richer paralinguistic features such as emotion and prosody, and generate speech responses directly within the backbone LLM rather than using a separate speech decoder. This integration also results in lower response latency and smoother interaction. However, native MLLMs suffer from catastrophic forgetting and performance degradation because the available paired speech-text data is insufficient to support the pretraining of MLLMs compared to the vast amount of text data required to pretrain text LLMs. To address this issue, we propose DeepTalk, a framework for adaptive modality expert learning based on a Mixture of Experts (MoE) architecture. DeepTalk first adaptively distinguishes modality experts according to their modality load within the LLM. Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training. As a result, DeepTalk incurs only a $5.5\%$ performance drop compared to the original LLM, which is significantly lower than the average performance drop of over $20\%$ typically seen in native MLLMs (such as GLM-4-Voice), and is on par with modular MLLMs. Meanwhile, the end-to-end dialogue latency remains within $0.5$ seconds, ensuring a seamless and intelligent speech interaction experience. Code and models are released at https://github.com/talkking/DeepTalk.
@@ 1 Introduction
Since GPT-4o GPT4o has demonstrated the great potential of using a unified model to process speech, recent voice interaction systems have evolved from traditional cascaded systems to end-to-end large speech interaction models. As shown in Figure 1 a), in traditional cascaded systems input speech is first transcribed into text by an automatic speech recognition (ASR) radford2023robust ; bai2024seed ; povey2011kaldi , the resulting text is then comprehended by a large language model (LLM) liu2024deepseek ; yang2024qwen2 ; glm2024chatglm to generate a text response, and finally the response is synthesized into speech by a text-to-speech (TTS) du2024cosyvoice ; ren2019fastspeech ; wang2025spark model. However, this cascaded structure results in high latency, and errors from each module can accumulate and propagate, leading to degraded performance. Therefore, recent speech LLMs adopt an end-to-end approach, using a unified model to process both speech and text simultaneously, which enhances the capabilities of speech understanding and generation. At the same time, this approach significantly reduces interaction latency and provides a more natural and smooth voice interaction experience.
@@
Current speech LLMs can be divided into two main categories: modular aligned multimodal and native multimodal chen2025minmo . Representatives of the former include Qwen2.5-Omni xu2025qwen2 , Minmo chen2025minmo , LLaMA-Omni fang2024llama , Freeze-Omni wang2024freeze , and VITA-1.5 fu2025vita . These models connect a speech encoder and a speech decoder to the LLM to handle audio understanding and audio generation tasks separately, placing semantic understanding within the LLM module. This maximally the retains LLM’s general capabilities. Furthermore, since the LLM itself does not require retraining, much less speech-text paired pretraining data is needed.
@@
Representative of the latter including Mini-Omni xie2024mini , GLM-4-Voice zeng2024glm , Moshi defossez2024moshi , and LUCY gao2025lucy use large amounts of audio-text paired data to retrain the LLM into a spoken language model, enabling simultaneous output of both audio and text tokens. Compared to modular aligned multimodal approaches, native ones greatly reduces the latency of speech token output. More paralinguistic information in speech, such as emotion and prosody, is retained within the SLM rather than in the speech decoder, benefiting the full expression of paralinguistic information. However, since native multimodal models expand the original LLM’s vocabulary and require retraining, they need a large amount of audio-text paired pretraining data. Therefore, in real-world scenarios where audio data is much less abundant than text data, native multimodal models are prone to catastrophic forgetting, severely impairing the original LLM’s language capabilities.
@@
In this paper, to alleviate the catastrophic forgetting faced by native multimodal models zhai2023investigating , we propose DeepTalk, a model that applies the mixture of experts (MoE) for multimodal learning li2025uni ; li2024cumo ; zhong2024moextend ; shen2024mome to effectively isolate modality-specific knowledge. Since most speech dialogue data are conversational and colloquial, while text LLMs are primarily trained on formal written text, training LLMs with colloquial speech dialogue data can lead to less standardized written outputs, resulting in decreased language performance on certain LLM benchmarks. Therefore, we assign different parameters within the LLM to specialize in instruction data from different modalities, allowing modality experts to focus on single-modality training and avoid interference between modalities. Finally, we use cross-modal instruction data to jointly train the modality experts, enabling the model to output both speech and text simultaneously.
@@
To minimize the damage to the original text LLM, we propose an adaptive modality expert selection strategy, which dynamically selects modality experts based on the MoE model’s modality load on different data. Experts with a high speech token load but a low text token load are selected as audio experts. After several iterations of modality expert selection and training, DeepTalk achieves only a loss of $5.5\%$ relative in language ability, reaching the same level of performance loss as modular multimodal models.
@@
  • We propose DeepTalk, which designates modality experts within the MoE to separate interference between modalities and isolate modality-specific knowledge. At the same time, modalities can collaborate to jointly output multimodal results, enabling both colloquial and formal language to coexist within a single model.
@@
Figure 1: Comparison of DeepTalk and existing voice interaction systems. Compared to modular MLLMs, DeepTalk embeds acoustic experts into a single pre-trained LLM and integrates both speech and language decoding within a single LLM. Through adaptive modality expert selection and a three-stage training process for modality experts, DeepTalk effectively alleviates catastrophic forgetting in text-based LLMs while maintaining robust voice interaction capabilities.
@@ 2.1 End-to-End Speech Interaction System
Speech is among the most important signal in human communication, as it not only conveys semantic content but also carries various paralinguistic information such as emotion, speech rate, intonation, and timbre. As a result, speech interaction has become an increasingly important mode of human-computer interaction. Traditional speech interaction systems are typically composed of a pipeline of separately optimized ASR, LLM, and TTS modules. Such systems not only suffer from high interaction latency, but also accumulate errors from each module, leading to a less satisfactory overall user experience. Therefore, end-to-end solutions that use a unified model to process both speech and text simultaneously have become the mainstream approach ji2024wavchat .
@@
Some of the early attempts integrate ASR into LLM by connecting the ASR encoder directly to the LLM via an adapter, rather than directly feeding the ASR transcript into the LLM kong2020panns ; chu2024qwen2 ; chu2023qwen ; das2024speechverse . This approach not only provides the LLM with richer speech information than using pure text input, but also reduces the latency caused by ASR recognition. However, these models still require an additional TTS module to achieve end-to-end speech interaction. To provide rich speech information, LLaMA-Omni fang2024llama feed LLM’s text hidden states to a non-autoregressive (NAR) Transformer based TTS module to predict discrete speech tokens. Since speech token sequences are much longer than text sequences for the same content, the text hidden states are upsampled and a Connectionist Temporal Classification (CTC) loss graves2006connectionist is used to align the upsampled hidden states with the audio token outputs. Meanwhile, to reduce latency for real-time speech interaction, a predefined chunk size is set to enable the vocoder to synthesize speech in a streaming manner. Mini-Omni employs SNAC, an audio codec, to discretize continuous acoustic signal into $7$ codebooks of discrete tokens at a rate of $82$Hz and adopts a delayed decoding strategy copet2023simple ; peng2024voicecraft ; lyth2024natural to model $7$ streams of acoustic tokens to effectively reduces the number of steps of LLM decoding during inference. Minmo chen2025minmo integrates a language model with an autoregressive CosyVoice2 decoder du2024cosyvoice . The hidden states of the language model are fed into the CosyVoice2 decoder to autoregressively predict speech tokens, which are then passed to a vocoder for waveform synthesis.
@@
Fully end-to-end approaches integrate the TTS module into the LLM, allowing the LLM itself to generate both text and speech tokens, rather than relying on a separate speech decoder or TTS module for speech synthesis. Discrete speech tokens are typically generated by neural audio codecs defossez2022high ; ji2024wavtokenizer ; zhang2023speechtokenizer or self-supervised models lakhotia2021generative ; hsu2021hubert ; chen2022wavlm . Currently, there are two mainstream modeling paradigm for end-to-end speech systems. The first is interleaved audio-text modeling, as exemplified by GLM-4-Voice zeng2024glm and Spirit-LM nguyen2025spirit , where audio and text tokens are interleaved within a single sequence, and the model alternately predicts audio and text tokens. The second is parallel audio-text modeling, represented by models such as Mini-Omni xie2024mini , Moshi defossez2024moshi , Slam-Omni chen2024slam and LUCY gao2025lucy . In this paradigm, text and audio tokens can be generated in parallel: while generating text tokens, multiple LM heads are used to simultaneously predict multiple audio tokens. Compared to the interleaved approach, parallel modeling compresses speech tokens into shorter sequences, enabling the modeling of higher-bitrate speech tokens.
@@
Modular Speech Language Models  Modular large speech models are primarily built on an architecture where a LLM is directly connected to an audio encoder and decoder through adapters. Examples include Qwen2.5-Omni xu2025qwen2 , Minmo chen2025minmo , LLaMA-Omni fang2024llama , and Freeze-Omni wang2024freeze , as shown in Fig. 1 b). In this structure, the LLM is only responsible for textual decoding, while speech decoding is handled by the audio decoder. As a result, the language capabilities of the LLM are largely preserved. Additionally, Minmo employs Low-Rank Adaptation (LoRA) hu2022lora for fine-tuning the LLM, while Freeze-Omni further freezes the LLM to prevent any degradation of its language abilities. However, such modular speech-language models face several challenges. For instance, the transmission of speech information relies entirely on the audio decoder, with the LLM only responsible for conveying textual content. This means that the expression of paralinguistic information in speech is completely dependent on the speech decoder, resulting in underutilization of the LLM. Furthermore, there are additional issues such as low deployment efficiency and the complexity of balancing the capacities of the LLM and the SLM.
@@
Native Speech Language Models Native speech language models reconstruct the LLM into an SLM, enabling the LLM to output both text and speech tokens simultaneously, as shown in Fig. 1 c). Currently, most mainstream native multimodal approaches such as Mini-Omni xie2024mini , Moshi defossez2024moshi , GLM-4-Voice zeng2024glm , and LUCY gao2025lucy have the LLM handle both text and speech decoding. This approach offers low latency, easy deployment, and allows a unified model to process both audio and text decoding, which better aligns with end-to-end requirements. However, without large-scale speech-text paired data, retraining the LLM in this way can lead to catastrophic forgetting and performance degradation.
@@ 2.2 Multimodal Mixture-of-Experts
BEiT-3 wang2023image , VLMo bao2022vlmo , and Uni-MoE li2025uni employ specific modality expert groups to capture modality-specific information. MoExtend zhong2024moextend further expands modality experts horizontally on top of existing LLMs to capture additional modality information. CuMo li2024cumo extends dense models into multimodal LLMs by co-upcycling the MLP, while MoME shen2024mome uses multiple feature encoders as modality expert groups to encode multi-dimensional features. VL-MoE shen2023scaling and MoE-LLaVa lin2024moe introduce mixture-of-experts (MoE) to improve training and deployment. MoMa lin2024moma pretrains MLLMs with multimodal mixture-of-experts and collaborates with sparse components to enhance the efficiency of training from scratch with trillions of mixed-modality tokens. Inspired by these works, we introduce multimodal mixture-of-experts (speech experts and text experts) into end-to-end large speech interaction models for native MLLM pretraining. At the same time, we use the modality load of each expert to adaptively select modality experts, a method called Adaptive Modality-Specific MoE, to address the catastrophic forgetting problem aforementioned.
@@ 3.1 Architecture
The overall architecture of the model is shown in Fig. 2, which mainly consists of an Audio Encoder, a connector, an MoE backbone, and a streaming codec. Let $X_{i}^{A}$ and $X_{i}^{T}$denote the audio input and text input of the $i$-th sample, respectively, and $Y_{i}^{A}$ and $Y_{i}^{T}$ denote the audio output and text output of the $i$-th sample, respectively. The model’s text and speech responses are represented as $Y^{T}\in V$ and $Y^{A}\in U$, where $V$ and $U$ denote the text vocabulary and the speech codec vocabulary, respectively. The embeddings of speech $H_{i}^{A}$ and text $H_{i}^{T}$ are combined and used as the overall feature $H_{i}$ input to the model, as shown below:
$H_{i}^{A}=Adapter(Encoder(X_{i}^{A})),$ (1)
$H_{i}^{T}=Embedding(X_{i}^{T}),$ (2)
$H_{i}=mean(H_{i}^{A},H_{i}^{T}),$ (3)
where $mean$ here refers to taking the average of the two embeddings. The loss L of the model over N samples can be defined as:
@@
where $T_{i}$ donates the maximum number of tokens of the output text $Y_{i}^{T}$ and the output speech $Y_{i}^{A}$ of the $i$-th sample.
@@
Audio Encoder and Adapter We use Whisper-medium radford2023robust as the audio encoder, which consists of 24 Transformer blocks with a hidden size of 4096 and 16 attention heads. It includes multiple convolutional downsampling layers that downsample the speech features by a factor of 4. We use an 80-dimensional log Mel-filterbank with a 25ms window length computed every 10ms as the input to the audio encoder. A convolution-based audio adapter is used to align the audio modality with the backbone language model and further downsample the speech features. Finally, for the input $X_{i}^{A}$, as shown in Equation 1, we obtain the hidden state $H_{i}^{A}$.
Figure 2: The DeepTalk model architecture.
@@
Table 1: Performance of Text to Speech on Seed-TTS. Lower values are better.
Model Seed-TTS $\downarrow$ $\downarrow$ $\downarrow$ Seed-TTS 1.12 2.25 7.59 CosyVoice 3.63 4.29 11.75 CosyVoice2 1.45 2.57 6.83 GLM-4-Voice 2.91 2.10 - VITA-1.5 8.44 2.63 - DeepTalk(topk=10) 1.87 7.34 7.53 DeepTalk(topk=1) 1.68 6.28 7.34 +DPO 1.41 3.25 7.29
@@
Mixture-of-Experts Backbone The backbone model is DeepSeek-V2-Lite liu2024deepseek , as shown in Fig. 2. Each MoE layer is adaptively divided into modality expert groups for specialized training and learning of the corresponding modality. Specifically, the text expert group is responsible for generating written text such as mathematical code, while the speech expert group is responsible for generating spoken language such as emotional speech response. In addition, there is a shared modality expert group for learning common knowledge across both modalities such as daily conversations. The weights among modality expert groups are allocated through modality routing, as follows:
$P(X_{l}^{AT})_{i}=\frac{e^{f(X_{l}^{AT})}}{\sum_{j}^{M+N}e^{f(X_{l}^{AT})_{j}}},$ (5)
$MoE(X_{l}^{AT})=\sum_{i=1}^{L}P(X_{l}^{AT})_{i}\cdot e(X_{l}^{AT})_{i},$ (6)
@@
where $X_{l}^{AT}$ denotes the averaged feature input of speech and text at the $l$-th layer, $L$ represents the total length of the sequence after concatenating speech and text, and $M$ and $N$ denote the number of speech experts and text experts, respectively. $f(x)=W\cdot x$ is the linear modality router to produce expert assignment probabilities and $W\in R^{h\times d}$. $h$ is the last dimension of hidden states and $d$ is the total number of experts.
@@
Audio-Text Modeling Following parallel modeling paradigm defossez2024moshi ; xie2024mini ; chen2024slam ; gao2025lucy , we use different heads to process hidden states, generating both text and audio tokens chen2024slam ; defossez2024moshi in one decoding step. Since the input to the LLM is the average of the text and audio representations instead of native text representations, maintaining the original capabilities of the LLM can be challenging. To enhance models’ text ability, we apply batch parallel decoding xie2024mini ; gao2025lucy , Batch-parallel decoding expands a single input to a batch size of two: the first sample generates text-only response and the second sample generates both text and speech response. The generated text tokens from the first sample are of better quality than those from the second and are used to substitute the latter in the second sample.
@@
Audio Codec Decoder We use SNAC codec siuzdak2024snac to encode the output speech into discrete tokens using $7$ codebooks, which have a total token rate of $82$ Hz. Combining speech and text tokens results in eight layers of labels, so we use eight LM heads to simultaneously predict one text token and seven audio token at each decoding step. We follow the delay pattern of MusicGen (copet2024simple, ) for better generation quality, applying $k$ token delay to $k$-th layer of audio tokens.

### v3

# DeepOmni: Towards Seamless and Smart Speech Interaction with Adaptive Modality-Specific MoE

@@ Abstract
Native multimodal large language models (MLLMs) restructure a single large language model (LLM) into a spoken language model (SLM) capable of both speech and text generation. Compared to modular and aligned MLLMs, native MLLMs preserve richer paralinguistic features such as emotion and prosody, and generate speech responses directly within the backbone LLM rather than using a separate speech decoder. This integration also results in lower response latency and smoother interaction. However, native MLLMs suffer from catastrophic forgetting and performance degradation because the available paired speech-text data is insufficient to support the pretraining of MLLMs compared to the vast amount of text data required to pretrain text LLMs. To address this issue, we propose DeepOmni, a framework for adaptive modality expert learning based on a Mixture of Experts (MoE) architecture. DeepOmni first adaptively distinguishes modality experts according to their modality load within the LLM. Each modality expert then undergoes specialized single-modality training, followed by joint multimodal collaborative training. As a result, DeepOmni incurs only a $5.5\%$ performance drop compared to the original LLM, which is significantly lower than the average performance drop of over $20\%$ typically seen in native MLLMs (such as GLM-4-Voice), and is on par with modular MLLMs. Meanwhile, the end-to-end dialogue latency remains within $0.5$ seconds, ensuring a seamless and intelligent speech interaction experience.
@@ 1 Introduction
Since GPT-4o OpenAI (2024) has demonstrated the great potential of using a unified model to process speech, recent voice interaction systems have evolved from traditional cascaded systems to end-to-end large speech interaction models. As shown in Figure 1 a), in traditional cascaded systems input speech is first transcribed into text by an automatic speech recognition (ASR) Radford et al. (2023); Bai et al. (2024); Povey et al. (2011), the resulting text is then comprehended by a large language model (LLM) Liu et al. (2024); Yang et al. (2024); GLM et al. (2024) to generate a text response, and finally the response is synthesized into speech by a text-to-speech (TTS) Du et al. (2024); Ren et al. (2019); Wang et al. (2025) model. However, this cascaded structure results in high latency, and errors from each module can accumulate and propagate, leading to degraded performance. Therefore, recent speech LLMs adopt an end-to-end approach, using a unified model to process both speech and text simultaneously, which enhances the capabilities of speech understanding and generation. At the same time, this approach significantly reduces interaction latency and provides a more natural and smooth voice interaction experience.
@@
Current speech LLMs can be divided into two main categories: modular aligned multimodal and native multimodal Chen et al. (2025). Representatives of the former include Qwen2.5-Omni Xu et al. (2025), Minmo Chen et al. (2025), LLaMA-Omni Fang et al. (2024), Freeze-Omni Wang et al. (2024), and VITA-1.5 Fu et al. (2025). These models connect a speech encoder and a speech decoder to the LLM to handle audio understanding and audio generation tasks separately, placing semantic understanding within the LLM module. This maximally the retains LLM’s general capabilities. Furthermore, since the LLM itself does not require retraining, much less speech-text paired pretraining data is needed.
@@
Representative of the latter including Mini-Omni Xie & Wu (2024), GLM-4-Voice Zeng et al. (2024), Moshi Défossez et al. (2024), LUCY Gao et al. (2025) and Step-Audio2 Wu et al. (2025) use large amounts of audio-text paired data to retrain the LLM into a spoken language model, enabling simultaneous output of both audio and text tokens. Compared to modular aligned multimodal approaches, native ones greatly reduces the latency of speech token output. More paralinguistic information in speech, such as emotion and prosody, is retained within the SLM rather than in the speech decoder, benefiting the full expression of paralinguistic information. However, since native multimodal models expand the original LLM’s vocabulary and require retraining, they need a large amount of audio-text paired pretraining data. Therefore, in real-world scenarios where audio data is much less abundant than text data, native multimodal models are prone to catastrophic forgetting, severely impairing the original LLM’s language capabilities.
@@
In this paper, to alleviate the catastrophic forgetting faced by native multimodal models Zhai et al. (2023), we propose DeepOmni, a model that applies the mixture of experts (MoE) for multimodal learning Li et al. (2025b; 2024); Zhong et al. (2024); Shen et al. (2024) to effectively isolate modality-specific knowledge. Since most speech dialogue data are conversational and colloquial, while text LLMs are primarily trained on formal written text, training LLMs with colloquial speech dialogue data can lead to less standardized written outputs, resulting in decreased language performance on certain LLM benchmarks. Therefore, we assign different parameters within the LLM to specialize in instruction data from different modalities, allowing modality experts to focus on single-modality training and avoid interference between modalities. Finally, we use cross-modal instruction data to jointly train the modality experts, enabling the model to output both speech and text simultaneously.
@@
To minimize the damage to the original text LLM, we propose an adaptive modality expert selection strategy, which dynamically selects modality experts based on the MoE model’s modality load on different data. Experts with a high speech token load but a low text token load are selected as audio experts. After several iterations of modality expert selection and training, DeepOmni achieves only a loss of $5.5\%$ relative in language ability, reaching the same level of performance loss as modular multimodal models.
@@
  • We propose DeepOmni, which designates modality experts within the MoE to separate interference between modalities and isolate modality-specific knowledge. At the same time, modalities can collaborate to jointly output multimodal results, enabling both colloquial and formal language to coexist within a single model.
@@
Figure 1: Comparison of DeepOmni and existing voice interaction systems. Compared to modular MLLMs, DeepOmni embeds acoustic experts into a single pre-trained LLM and integrates both speech and language decoding within a single LLM. Through adaptive modality expert selection and a three-stage training process for modality experts, DeepOmni effectively alleviates catastrophic forgetting in text-based LLMs while maintaining robust voice interaction capabilities.
@@ 2.1 End-to-End Speech Interaction System
Speech is among the most important signal in human communication, as it not only conveys semantic content but also carries various paralinguistic information such as emotion, speech rate, intonation, and timbre. As a result, speech interaction has become an increasingly important mode of human-computer interaction. Traditional speech interaction systems are typically composed of a pipeline of separately optimized ASR, LLM, and TTS modules. Such systems not only suffer from high interaction latency, but also accumulate errors from each module, leading to a less satisfactory overall user experience. Therefore, end-to-end solutions that use a unified model to process both speech and text simultaneously have become the mainstream approach Ji et al. (2024a).
@@
Some of the early attempts integrate ASR into LLM by connecting the ASR encoder directly to the LLM via an adapter, rather than directly feeding the ASR transcript into the LLM Kong et al. (2020); Chu et al. (2024; 2023); Das et al. (2024). This approach not only provides the LLM with richer speech information than using pure text input, but also reduces the latency caused by ASR recognition. However, these models still require an additional TTS module to achieve end-to-end speech interaction. To provide rich speech information, LLaMA-Omni Fang et al. (2024) feed LLM’s text hidden states to a non-autoregressive (NAR) Transformer based TTS module to predict discrete speech tokens. Since speech token sequences are much longer than text sequences for the same content, the text hidden states are upsampled and a Connectionist Temporal Classification (CTC) loss Graves et al. (2006) is used to align the upsampled hidden states with the audio token outputs. Meanwhile, to reduce latency for real-time speech interaction, a predefined chunk size is set to enable the vocoder to synthesize speech in a streaming manner. Mini-Omni employs SNAC, an audio codec, to discretize continuous acoustic signal into $7$ codebooks of discrete tokens at a rate of 8282Hz and adopts a delayed decoding strategy Copet et al. (2023); Peng et al. (2024); Lyth & King (2024) to model $7$ streams of acoustic tokens to effectively reduces the number of steps of LLM decoding during inference. Minmo Chen et al. (2025) integrates a language model with an autoregressive CosyVoice2 decoder Du et al. (2024). The hidden states of the language model are fed into the CosyVoice2 decoder to autoregressively predict speech tokens, which are then passed to a vocoder for waveform synthesis.
@@
Fully end-to-end approaches integrate the TTS module into the LLM, allowing the LLM itself to generate both text and speech tokens, rather than relying on a separate speech decoder or TTS module for speech synthesis. Discrete speech tokens are typically generated by neural audio codecs Défossez et al. (2022); Ji et al. (2024b); Zhang et al. (2023) or self-supervised models Lakhotia et al. (2021); Hsu et al. (2021); Chen et al. (2022). Currently, there are two mainstream modeling paradigm for end-to-end speech systems. The first is interleaved audio-text modeling, as exemplified by GLM-4-Voice Zeng et al. (2024), Baichuan-Audio Li et al. (2025a) and Spirit-LM Nguyen et al. (2025), where audio and text tokens are interleaved within a single sequence, and the model alternately predicts audio and text tokens. The second is parallel audio-text modeling, represented by models such as Mini-Omni Xie & Wu (2024), Moshi Défossez et al. (2024), Slam-Omni Chen et al. (2024) and LUCY Gao et al. (2025). In this paradigm, text and audio tokens can be generated in parallel: while generating text tokens, multiple LM heads are used to simultaneously predict multiple audio tokens. Compared to the interleaved approach, parallel modeling compresses speech tokens into shorter sequences, enabling the modeling of higher-bitrate speech tokens.
@@
Modular Speech Language Models  Modular large speech models are primarily built on an architecture where a LLM is directly connected to an audio encoder and decoder through adapters. Examples include Qwen2.5-Omni Xu et al. (2025), Minmo Chen et al. (2025), LLaMA-Omni Fang et al. (2024), and Freeze-Omni Wang et al. (2024), as shown in Fig. 1 b). In this structure, the LLM is only responsible for textual decoding, while speech decoding is handled by the audio decoder. As a result, the language capabilities of the LLM are largely preserved. Additionally, Minmo employs Low-Rank Adaptation (LoRA) Hu et al. (2022) for fine-tuning the LLM, while Freeze-Omni further freezes the LLM to prevent any degradation of its language abilities. However, such modular speech-language models face several challenges. For instance, the transmission of speech information relies entirely on the audio decoder, with the LLM only responsible for conveying textual content. This means that the expression of paralinguistic information in speech is completely dependent on the speech decoder, resulting in underutilization of the LLM. Furthermore, there are additional issues such as low deployment efficiency. In modular speech language models paradigm, the LLM is trained to model only text distribution, or formally $P(T_{out}|T_{in}\cup S_{in})$, and the audio decoder is trained to model $P(S_{out}|T_{out})$.
@@
Native Speech Language Models Native speech language models reconstruct the LLM into an SLM, enabling the LLM to output both text and speech tokens simultaneously, as shown in Fig. 1 c). Currently, most mainstream native multimodal approaches such as Mini-Omni Xie & Wu (2024), Moshi Défossez et al. (2024), GLM-4-Voice Zeng et al. (2024), and LUCY Gao et al. (2025) have the LLM handle both text and speech decoding. This approach offers low latency, easy deployment, and allows a unified model to process both audio and text decoding, which better aligns with end-to-end requirements. However, without large-scale speech-text paired data, retraining the LLM in this way can lead to catastrophic forgetting and performance degradation. In native speech language models paradigm, the LLM is trained to model $P(T_{out}\cup S_{out}|T_{in}\cup S_{in})$ . Clearly in the native paradigm, the speech-text joint distribution space for the LLM to learn is more complex than text distribution. It is also for this reason in modular paradigm the LLM does not need to be restructed to a SLM and its language capacity is largely preserved.
@@ 2.2 Multimodal Mixture-of-Experts
Multimodal Mixture-of-Experts refers to using a MoE model to learn multimodal knowledge. BEiT-3 Wang et al. (2023), VLMo Bao et al. (2022), and Uni-MoE Li et al. (2025b) employ specific modality expert groups to capture modality-specific information. MoExtend Zhong et al. (2024) further expands modality experts horizontally on top of existing LLMs to capture additional modality information. CuMo Li et al. (2024) extends dense models into multimodal LLMs by co-upcycling the MLP, while MoME Shen et al. (2024) uses multiple feature encoders as modality expert groups to encode multi-dimensional features. VL-MoE Shen et al. (2023) and MoE-LLaVa Lin et al. (2024a) introduce mixture-of-experts (MoE) to improve training and deployment. MoMa Lin et al. (2024b) pretrains MLLMs with multimodal mixture-of-experts and collaborates with sparse components to enhance the efficiency of training from scratch with trillions of mixed-modality tokens. Inspired by these works, we introduce multimodal mixture-of-experts (speech experts and text experts) into end-to-end large speech interaction models for native MLLM pretraining. At the same time, we use the modality load of each expert to adaptively select modality experts, a method called Adaptive Modality-Specific MoE, to address the catastrophic forgetting problem aforementioned.
@@ 3.1 Architecture
The overall architecture of the model is shown in Fig. 2, which mainly consists of an Audio Encoder, a connector, an MoE backbone, and a streaming codec. Let $X_{i}^{A}$ and $X_{i}^{T}$denote the audio input and text input of the $i$-th sample, respectively, and $Y_{i}^{A}$ and $Y_{i}^{T}$ denote the audio output and text output of the $i$-th sample, respectively. The model’s text and speech responses are represented as $Y^{T}\in V$ and $Y^{A}\in U$, where $V$ and $U$ denote the text vocabulary and the speech codec vocabulary, respectively. The embeddings of speech $H_{i}^{A}$ and text $H_{i}^{T}$ are combined and used as the overall feature $H_{i}$ input to the model. The loss L of the model over N samples can be defined as:
@@
where $T_{i}$ denotes the maximum number of tokens of the output text $Y_{i}^{T}$ and the output speech $Y_{i}^{A}$ of the $i$-th sample.
@@
Audio Encoder and Adapter We use Whisper-medium Radford et al. (2023) as the audio encoder, which consists of 24 Transformer blocks with a hidden size of 4096 and 16 attention heads. It includes multiple convolutional downsampling layers that downsample the speech features by a factor of 4. We use an 80-dimensional log Mel-filterbank with a 25ms window length computed every 10ms as the input to the audio encoder. A MLP-based audio adapter is used to align the audio modality with the backbone language model and further downsample the speech features. Finally, for the input $X_{i}^{A}$, we obtain the hidden state $H_{i}^{A}=Adapter(Encoder(X_{i}^{A}))$.
Figure 2: The DeepOmni model architecture.
@@
Table 1: Performance of Text to Speech on Seed-TTS. Lower values are better. The metric for Chinese is character error rate (CER), and the metric for English is word error rate (WER).
[…]
@@
Mixture-of-Experts Backbone The backbone model is DeepSeek-V2-Lite Liu et al. (2024), as shown in Fig. 2. Each MoE layer is adaptively divided into modality expert groups for specialized training and learning of the corresponding modality. Specifically, the text expert group is responsible for generating written text such as mathematical code, while the speech expert group is responsible for generating spoken language such as emotional speech response. In addition, there is a shared modality expert group for learning common knowledge across both modalities such as daily conversations. The weights among modality expert groups are allocated through modality routing, as follows:
$P(h_{l})_{i}=\frac{e^{f(h_{l})_{i}}}{\sum_{j}^{M}e^{f(h_{l})_{i,j}}},$ (2)
$MoE(h_{l})=\sum_{i=1}^{T}P(h_{l})_{i}\cdot e(h_{l})_{i},$ (3)
@@
where $h_{l}$ denotes the averaged feature input of speech and text at the $l$-th layer, $T$ represents the total length of the sequence after concatenating speech and text, and $M$ denotes the total number of speech and text experts combined, respectively. $f(x)=W\cdot x$ is the linear modality router to produce expert assignment probabilities and $W\in R^{h\times d}$. $h$ is the last dimension of hidden states and $d$ is the total number of experts.
@@
Audio-Text Modeling Following parallel modeling paradigm Défossez et al. (2024); Xie & Wu (2024); Chen et al. (2024); Gao et al. (2025), we use different heads to process hidden states, generating both text and audio tokens Chen et al. (2024); Défossez et al. (2024) in one decoding step. Since the input to the LLM is the average of the text and audio representations instead of native text representations, maintaining the original capabilities of the LLM can be challenging. To enhance models’ text ability, we apply batch parallel decoding Xie & Wu (2024); Gao et al. (2025), Batch-parallel decoding expands a single input to a batch size of two: the first sample generates text-only response and the second sample generates both text and speech response. The generated text tokens from the first sample are of better quality than those from the second and are used to substitute the latter in the second sample.
@@
Audio Codec Decoder We use SNAC codec Siuzdak et al. (2024) to encode the output speech into discrete tokens using $7$ codebooks, which have a total token rate of $82$ Hz. Combining speech and text tokens results in eight layers of labels, so we use eight LM heads to simultaneously predict one text token and seven audio token at each decoding step. We follow the delay pattern of MusicGen (Copet et al., 2024) for better generation quality, applying $k$ token delay to $k$-th layer of audio tokens.